# Many Heads Are Better Than One: Improved Scientific Idea Generation by A LLM-Based Multi-Agent System

### v2

@@ 1 Introduction
The rapid acceleration of scientific research necessitates innovative tools for exploring new concepts and tackling complex challenges (park2023papers, ). Automatic scientific discovery has emerged as a promising solution to accelerate innovation, aligning with the ultimate goal within the scientific community (langley1987scientific, ). With the development of artificial intelligence (AI), automatic scientific discovery has the potential to revolutionize how research is conducted by automating key steps in the scientific process, ranging from hypothesis generation to experimental design (raghu2020survey, ; spangler2014automated, ).
Recent works like AI Scientist (lu2024ai, ), ResearchTown (yu2024researchtown, ), and HypoGen qi2024large leverage LLMs (openai, ; llama3, ) as intelligent agents to simulate the scientific idea generation process and advance automatic scientific discovery at various stages, including literature review (shi2023towards, ; hsu2024chime, ) and experimental design (wu2023autogen, ; huang2024crispr, ). However, these efforts either rely on a single-agent system, overlooking the collaborative nature of real-world research (kayacik2019identifying, ; gauch2003scientific, ; linsey2005collaborating, ), or employ an oversimplified collaboration framework and unrealistic data (e.g. personal information, collaboration networks) to model a multi-agent system, failing to capture real-world complexities. Consequently, they provide limited insights into multi-agent collaboration, which is essential for advancing autonomous scientific discovery.
@@
To address these limitations, we build a virtual Scientific Research Ecosystem to benchmark the potential of multi-agent systems in scientific idea generation. Specifically, Scientific Research Ecosystem is a digit twin of research communities at the given time point, e.g., Jan 1st, 2024. It have three components: (1) 35,774,510 virtual scientists whose backgrounds and publications are cloned from real scientists at that moment, (2) a past paper database where papers are published before that moment and (3) a contemporary paper database where papers are published after that moment. To evaluate the novelty of the generated ideas, we introduce three metrics from different perspectives: dissimilarity to past papers, alignment with research trends, and the potential influence on contemporary research (shao2020bert, ; yang2022gender, ). By comparing the generated abstract against two paper databases, we ensure the generated ideas are innovative and align with emerging scientific directions, validating the effectiveness of our approach (Real examples and their professional analysis are shown in Appx. D).
Based on the virtual Scientific Research Ecosystems, we propose an LLM-based multi-agent system, Virtual Scientists (VirSci), designed to harness the potential of LLM agents in assisting autonomous scientific idea generation. Leveraging the inherent human-like reasoning capabilities of LLMs (xie2024can, ), VirSci simulates the collaborative process of generating scientific ideas (perry2017creativity, ; muzzio2024collaborative, ), by the following five steps (see Fig. 1): (1) Collaborator Selection, (2) Topic Selection, (3) Idea Generation, (4) Idea Novelty Assessment, and (5) Abstract Generation. In Collaborator Selection stage, given a randomly selected agent as the team leader, it will exploit historical co-authors based on its collaboration history and academic social networks, while also exploring potential collaborators whose expertise and research interests align with the team’s goals march1991exploration ; rzhetsky2015choosing ; zeng2019increasing .
@@
In Topic Selection stage, the scientists will discuss topics of common interest. The discussion will be terminated and restarted if a consensus cannot be achieved among the majority of the team. The scientists who are not interested can choose to leave the discussion at will. Otherwise, the discussion continues until a final topic is determined. In the Idea Generation stage, the virtual scientists retrieve relevant papers from the past paper database and engage in both inter- and intra-team discussions, where collaborators participate in iterative dialogues based on their backgrounds. This inter- and intra-team discussion distinguishes our approach from previous group discussion patterns zhang2023exploring ; qi2024large ; qian2024chat , enabling agents within the team to proactively seek advice from external agents (inter-team) through an “Invitation Mechanism”, while effectively balancing diverse perspectives within the team (intra-team). Once the discussions are over, the Abstract Generation stage begins, and the team generates a comprehensive abstract that encapsulates the proposed ideas.
@@
We conduct extensive experiments to verify the effectiveness of VirSci in producing novel scientific ideas on both single-discipline and multi-discipline datasets. The findings prove that the multi-agent system improves the single-agent executive by, on average, +13.8% and +44.1% in alignment with and potential impacts on contemporary research, respectively. Furthermore, our experiments investigate collaboration mechanisms among agents that influence the performance of idea generation. The patterns observed in the experimental results align with findings from prior Science of Science studies (fortunato2018science, ; wu2019large, ; zeng2021fresh, ; shi2023surprising, ) published in Top venues, e.g., Science and Nature, providing valuable insights to guide future research toward autonomous scientific discovery.
Our core contributions are as follows:
@@ 2.1 AI for Scientific Discovery
In recent years, AI has transformed scientific discovery by offering powerful tools to enhance research processes (xu2021artificial, ). Generative AI, in particular, accelerates basic scientific discoveries by tackling complex tasks such as molecular identification (vignac2022digress, ), protein structure prediction (abramson2024accurate, ), and proteomics research (ding2024automating, ), significantly reducing experimental iteration times.
Besides, with the advent of LLMs, AI methodologies can step further and collaborate in streamlining critical stages of the scientific pipeline, including hypothesis generation, experimental design, data acquisition, and analysis(zheng2023large, ; wang2023scientific, ; miret2024llms, ; wysocki2024llm, ; lu2024ai, ; si2024can, ). Nevertheless, these approaches lack the collaborative nature of the scientists intrinsic to real-world research. VirSci is the first to harness the power of an LLM-based multi-agent system to facilitate the generation of research ideas in autonomous scientific discovery.
@@ 2.2 Collaboration in Multi-Agent Systems
A multi-agent system for team collaboration leverages autonomous agents to coordinate, communicate, and solve tasks within a shared environment, simulating human teamwork dynamics (dorri2018multi, ). Traditional systems typically rely on semi-autonomous agents using explicit protocols and structured messages to achieve shared goals (dunin2011teamwork, ; bakliwal2018multi, ). The emergence of LLMs has reshaped this landscape by enabling agents to communicate in natural language, fostering more intuitive and flexible interactions (park2023generative, ). Studies have demonstrated that LLM-based multi-agent systems outperform single-agent setups in tasks such as programming, game playing, and complex reasoning (du2024multi, ; wang2024rethinking, ; light2023avalonbench, ). However, prior multi-agent frameworks (qi2024large, ; yu2024researchtown, ) for focused scientific idea generation (1) lack dynamic team composition, restricting team members from accessing insights beyond their initial group, and (2) do not incorporate real scientist data and collaboration relationship, leading to impractical conclusions. To overcome these limitations, we introduce an inter- and intra-team discussion mechanism, enabling flexible collaboration among virtual scientists (supported by real data) and fostering the generation of de novo scientific ideas.
@@ 3.1 The Scientific Research Ecosystem
Past Paper Database. To construct the past paper database $B_{\text{past}}$ using the Faiss [^4] Faiss is a Python library designed for efficient similarity search and clustering of dense vectors. johnson2019billion , we selected papers published before the $y_{bound}$. Each paper includes essential information such as its title, citation count, and abstract.
@@
Author Knowledge Bank. For each scientist in $S$, we extract their basic profile from the pre-processed dataset (More details are shown in Sec. 4.1), which includes their name, affiliations, citation count, research interests, and collaboration history. Using the KnowledgeBank module from AgentScope (agentscope, ), we embed these scientist profiles into the author knowledge bank. This allows agents to quickly access and familiarize themselves with other initialized agents’ information. Notably, real author names are masked to prevent data leakage and privacy problems during agent initialization.
Adjacency Matrix. Given the scientist set $S$, let $A$ represent the adjacency matrix, where $A_{i,j}$ denotes the number of times that scientist $i$ has collaborated with scientist $j$. To prevent agents from always choosing previously collaborated scientists, overlooking the benefit of fresh collaborations that often lead to more original and impactful research (zeng2021fresh, ), we increment all values in $A$ by 1 (more experiments about the increment function in Sec. 5.4). This adjustment ensures an explore-exploit mechanism, allowing scientists with no prior collaborations a chance to be selected, thereby encouraging agents to explore new partnerships.
@@ 3.2 The Multi-agent System
We first randomly sample a scientist $s_{0}$ from $S$ as the team leader. The team leader then follows these steps to produce an abstract: (1) Collaborator Selection, (2) Topic Discussion, (3) Idea Generation, (4) Idea Novelty Assessment, and (5) Abstract Generation. To help each agent become familiar with the backgrounds of other team members without overloading the initialization prompt, we employ retrieval-augmented generation (RAG) (lewis2020retrieval, ), used throughout all five steps. All necessary prompts and example scenarios are shown in Appx. E and F.
Collaborator Selection. The first step in our system is to form a team of scientists, $T=\{s_{0},\dots,s_{i},\dots,s_{n}\}$, where $n$ denotes the team size. When $s_{0}$ is selecting collaborators, we convert the adjacency matrix, $A$, into a probability distribution using the following equation: $P_{i,j}=\frac{A_{i,j}}{\sum_{j=1}^{N}A_{i,j}}$, where $N$ denotes the size of $S$. This allows the team leader to iteratively send invitations to preferred collaborators. Upon receiving an invitation, the invited scientist evaluates whether to join the team using the chain-of-thought process (wei2022chain, ), considering the profiles of $s_{0}$ and the current team members. If accepted, the scientist is added to the team $T$. This process continues until the pre-defined team size $n$ is reached.
Topic Discussion. The next step is to propose a research topic, which will guide the research direction. Inspired by multi-round collaboration (mezirow2003critical, ; sunstein2005societies, ; amgoud2009using, ) and multi-agent collaboration strategies (xu2023magic, ; zhang2023exploring, ; shinn2024reflexion, ), we design an inter- and intra-team discussion mechanism. In this mechanism, team members engage in discussions based on a specific task description prompt (inter-team discussion). The default discussion mechanism between agents follows a round-table format with a sequential progression (Additional discussion topologies are presented in Sec. 5.3.1). This process is also applied to subsequent collaboration steps. While allowing agents to decide when to stop the discussion would better reflect real-world scenarios, fixing the number of turns ensures consistent inference costs across different team settings in our experiments. Therefore, we leave the discussion of adaptive turn numbers to the ablation study (See Sec. 5.3.2). The prompt for agent $i$ during the topic discussion is:
@@
Idea Generation. Third, the team is tasked with proposing several potential ideas. To align with genuine research workflows and mitigate LLM illusions (huang2023survey, ), each agent is required to generate a comprehensive response that includes three key components: (1) a description of the idea, (2) a specific experimental plan, and (3) a self-assessment covering metrics such as novelty, feasibility, and clarity, representing the agent’s confidence (See Appx. Fig. 20).
@@
Inspired by the concept of gradually expanding an archive of ideas (zhang2023exploring, ; lu2024ai, ), when a scientist $s_{i}$ at turn $k$ receives an existing idea from the previous response $R_{k,i-1}$, we retain the previously generated ideas along with their corresponding references from $B_{past}$. These are passed to the next agent, who can either refine the existing idea or propose a new one, depending on its choice. The prompt $Q_{k,i}$ is represented as:
@@
Afterwards, the response of $S_{i}$ at turn $k$ can be represented as $R_{k,i}\sim P_{s_{i}}(\cdot\mid Q_{k,i})$. After $K$ turns of discussion, we retain the three ideas with the highest confidence and store them in the idea list $I$.
@@
where $I_{j}$ is the $j$-th idea in $I$. Following chain-of-thought process, the response $R_{k,i}\sim P_{s_{i}}(\cdot\mid Q_{k,i})$ includes the scientist’s preferred idea and the reasoning behind their choice. The idea receiving the highest votes is then selected as the final idea, $R_{idea}$, for abstract generation. The effectiveness of novelty assessment is discussed in Sec. 5.2.
Abstract Generation. Lastly, the team is required to produce a comprehensive abstract that includes the following sections: (1) Introduction, (2) Objective, (3) Methods, (4) Expected Results, and (5) Conclusion (alexandrov2007writing, ). At the start of abstract generation, the team leader provides an initial draft based on $R_{idea}$. The first abstract-generation prompt is: $Q_{1,0}=\langle Q_{abstract},R_{idea}\rangle$, where $Q_{abstract}$ represents the task description and format requirements.
When an abstract is provided by the previous response $R_{k,i-1}$, the next scientist’s response should include: (1) an evaluation of the prior abstract (evaluation metrics are detailed in Appx. Fig. 23), (2) proposed modifications, and (3) the revised abstract to enable continuous refinement. The prompt is defined as:
@@
A self-review mechanism is also considered after $R_{abstract}$ is finalized to pre-check its novelty. In this self-review, the optimized abstract $R_{abstract}$ is provided to the team leader to assess novelty by comparing it to similar papers in $B_{past}$. The prompt is:
If the abstract undergoes a second self-review and still does not meet the novelty requirement, it will be discarded, and the team will generate a new idea. Once the self-review yields satisfactory results, the final abstract will be produced, and the system will terminate. However, this self-review mechanism introduces uncertainty in total inference cost, making it difficult to ensure fair experimental comparisons. We discuss the effectiveness of this module only in the ablation study (see Sec. 5.2).

### v3

@@ Abstract
Many Heads Are Better Than One: Improved Scientific Idea Generation
by A LLM-Based Multi-Agent System
Haoyang Su1, [^1] Equal contribution., Renqi Chen1, [^1] Equal contribution., Shixiang Tang1,5, [^2] Corresponding authors: Nanqing Dong (dongnanqing@pjlab.org.cn) and Shixiang Tang (tangshixiang@pjlab.org.cn)., Zhenfei Yin3, Xinzhe Zheng1, Jinzhe Li1, Biqing Qi1, Qi Wu4, Hui Li4, Wanli Ouyang1,5,Philip Torr3, Bowen Zhou1,6, Nanqing Dong1,2, [^2] Corresponding authors: Nanqing Dong (dongnanqing@pjlab.org.cn) and Shixiang Tang (tangshixiang@pjlab.org.cn)., [^3] Project lead. 1Shanghai Artificial Intelligence Laboratory 2Shanghai Innovation Institute 3Department of Engineering Science, University of Oxford 4Shanghai Institute for Science of Science 5Department of Information Engineering, Chinese University of Hong Kong 6Department of Electronic Engineering, Tsinghua University
@@ 1 Introduction
The rapid acceleration of scientific research necessitates innovative tools for exploring new concepts and tackling complex challenges (Park et al., 2023b). Automatic scientific discovery has emerged as a promising solution to accelerate innovation, aligning with the ultimate goal within the scientific community (Langley, 1987). With the development of artificial intelligence (AI), automatic scientific discovery has the potential to revolutionize how research is conducted by automating key steps in the scientific process, ranging from hypothesis generation to experimental design (Raghu and Schmidt, 2020; Spangler et al., 2014).
Recent works like AI Scientist (Lu et al., 2024), ResearchTown Yu et al. (2024), and HypoGen Qi et al. (2024) leverage LLMs (OpenAI, 2023; Dubey et al., 2024) as intelligent agents to simulate the scientific idea generation process and advance automatic scientific discovery at various stages, including literature review (Shi et al., 2023; Hsu et al., 2024) and experimental design (Wu et al., 2023; Huang et al., 2024). However, these efforts either rely on a single-agent system, overlooking the collaborative nature of real-world research (Kayacik et al., 2019; Gauch, 2003; Linsey et al., 2005), or employ an oversimplified collaboration framework and unrealistic data (e.g. manually crafted personal profiles, synthetic collaboration networks) to model a multi-agent system, failing to capture the dynamic relationships that characterize real scientific teams. Consequently, they provide limited insights into multi-agent collaboration, which is essential for advancing autonomous scientific discovery.
@@
To address these limitations, we build a virtual Scientific Research Ecosystem to benchmark the potential of multi-agent systems in scientific idea generation. Specifically, Scientific Research Ecosystem is a digit twin of research communities at the given time point, e.g., Jan 1st, 2024. It have three components: (1) virtual scientists whose backgrounds and publications are cloned from real scientists at that moment, (2) a past paper database where papers are published before that moment and (3) a contemporary paper database where papers are published after that moment. To evaluate the novelty of the generated ideas, we introduce three metrics from different perspectives: dissimilarity to past papers, alignment with research trends, and the potential influence on contemporary research (Shao et al., 2020; Yang et al., 2022). By comparing the generated abstract against two paper databases, we ensure the generated ideas are innovative and align with emerging scientific directions, validating the effectiveness of our approach. Real examples and their professional analysis are shown in Appx. F.
@@
Based on the virtual Scientific Research Ecosystems, we propose an LLM-based multi-agent system, Virtual Scientists (VirSci), designed to harness the potential of LLM agents in assisting autonomous scientific idea generation. Leveraging the inherent human-like reasoning capabilities of LLMs (Xie et al., 2024), VirSci simulates the collaborative process of generating scientific ideas (Perry-Smith and Mannucci, 2017; Muzzio and Gama, 2024), by the following five steps (see Fig. 1): (1) Collaborator Selection, (2) Topic Selection, (3) Idea Generation, (4) Idea Novelty Assessment, and (5) Abstract Generation. In Collaborator Selection stage, given a randomly selected agent as the team leader, it will exploit historical co-authors based on its collaboration history and academic social networks, while also exploring potential collaborators whose expertise and research interests align with the team’s goals March (1991); Rzhetsky et al. (2015); Zeng et al. (2019). In Topic Selection stage, the scientists will discuss topics of common interest. The discussion will be terminated and restarted if a consensus cannot be achieved among the majority of the team. The scientists who are not interested can choose to leave the discussion at will. Otherwise, the discussion continues until a final topic is determined. In the Idea Generation stage, the virtual scientists retrieve relevant papers from the past paper database and engage in both inter- and intra-team discussions, where collaborators participate in iterative dialogues based on their backgrounds. This inter- and intra-team discussion distinguishes our approach from previous group discussion patterns Zhang et al. (2024); Qi et al. (2024); Qian et al. (2024), enabling agents within the team to proactively seek advice from external agents (inter-team) through an “Invitation Mechanism”, while effectively balancing diverse perspectives within the team (intra-team). Once the discussions are over, the Abstract Generation stage begins, and the team generates a comprehensive abstract that encapsulates the proposed ideas.
@@
We conduct extensive experiments to verify the effectiveness of VirSci in producing novel scientific ideas on both single-discipline and multi-discipline datasets. The findings prove that the multi-agent system improves the single-agent executive by, on average, +13.8% and +44.1% in alignment with and potential impacts on contemporary research, respectively. Furthermore, our experiments investigate collaboration mechanisms among agents that influence the performance of idea generation. The patterns observed in the experimental results align with findings from prior Science of Science studies (Fortunato et al., 2018; Wu et al., 2019; Zeng et al., 2021; Shi and Evans, 2023) published in Top venues, e.g., Science and Nature, providing valuable insights to guide future research toward autonomous scientific discovery. Our core contributions are as follows:
@@ 2.1 AI for Scientific Discovery
In recent years, AI has transformed scientific discovery by offering powerful tools to enhance research processes (Xu et al., 2021). Generative AI, in particular, accelerates basic scientific discoveries by tackling complex tasks such as molecular identification (Vignac et al., 2022), protein structure prediction (Abramson et al., 2024), and proteomics research (Ding et al., 2024), significantly reducing experimental iteration times. Besides, with the advent of LLMs, AI methodologies can step further and collaborate in streamlining critical stages of the scientific pipeline, including hypothesis generation, experimental design, data acquisition, and analysis(Zheng et al., 2023; Wang et al., 2023; Miret and Krishnan, 2024; Wysocki et al., 2024; Lu et al., 2024; Si et al., 2024). Nevertheless, these approaches lack the collaborative nature of the scientists intrinsic to real-world research. VirSci is the first to harness the power of an LLM-based multi-agent system to facilitate the generation of research ideas in autonomous scientific discovery.
@@ 2.2 Collaboration in Multi-Agent Systems
A multi-agent system for team collaboration leverages autonomous agents to coordinate, communicate, and solve tasks within a shared environment, simulating human teamwork dynamics (Dorri et al., 2018). Traditional systems typically rely on semi-autonomous agents using explicit protocols and structured messages to achieve shared goals (Dunin-Keplicz and Verbrugge, 2011; Bakliwal et al., 2018). The emergence of LLMs has reshaped this landscape by enabling agents to communicate in natural language, fostering more intuitive and flexible interactions (Park et al., 2023a). Studies have demonstrated that LLM-based multi-agent systems outperform single-agent setups in tasks such as programming, game playing, and complex reasoning (Du et al., 2024; Wang et al., 2024; Light et al., 2023). However, prior multi-agent frameworks (Qi et al., 2024; Yu et al., 2024) for focused scientific idea generation (1) lack dynamic team composition, restricting team members from accessing insights beyond their initial group, and (2) do not incorporate real scientist data and collaboration relationship, leading to impractical conclusions. To overcome these limitations, we introduce an inter- and intra-team discussion mechanism, enabling flexible collaboration among virtual scientists (supported by real data) and fostering the generation of de novo scientific ideas.
@@ 3.1 The Scientific Research Ecosystem
Past Paper Database. To construct the past paper database $B_{\text{past}}$ using the Faiss [^2] Faiss is a Python library designed for efficient similarity search and clustering of dense vectors. Johnson et al. (2019), we selected papers published before the $y_{bound}$. Each paper includes essential information such as its title, citation count, and abstract.
@@
Author Knowledge Bank. For each scientist in $S$, we extract their basic profile from the pre-processed dataset (More details are shown in Sec. 4.1), which includes their name, affiliations, citation count, research interests, and collaboration history. Using the KnowledgeBank module from AgentScope (Gao et al., 2024), we embed these scientist profiles into the author knowledge bank. This allows agents to quickly access and familiarize themselves with other initialized agents’ information. Notably, real author names are masked to prevent data leakage and privacy problems during agent initialization (See Sec. 7).
Adjacency Matrix. Given the scientist set $S$, let $A$ represent the adjacency matrix, where $A_{i,j}$ denotes the number of times that scientist $i$ has collaborated with scientist $j$. To prevent agents from always choosing previously collaborated scientists, overlooking the benefit of fresh collaborations that often lead to more original and impactful research (Zeng et al., 2021), we increment all values in $A$ by 1 (more experiments about the increment function in Appx. D.4). This adjustment ensures an explore-exploit mechanism, allowing scientists with no prior collaborations a chance to be selected, thereby encouraging agents to explore new partnerships.
@@ 3.2 The Multi-agent System
We first randomly sample a scientist $s_{0}$ from $S$ as the team leader. The team leader then follows these steps to produce an abstract: (1) Collaborator Selection, (2) Topic Discussion, (3) Idea Generation, (4) Idea Novelty Assessment, and (5) Abstract Generation. To help each agent become familiar with the backgrounds of other team members without overloading the initialization prompt, we employ retrieval-augmented generation (RAG) (Lewis et al., 2020), used throughout all five steps. All necessary prompts and example scenarios are shown in Appx. H and I.
Collaborator Selection. The first step in our system is to form a team of scientists, $T=\{s_{0},\dots,s_{i},\dots,s_{n}\}$, where $n$ denotes the team size. When $s_{0}$ is selecting collaborators, we convert the adjacency matrix, $A$, into a probability distribution using the following equation: $P_{i,j}=\frac{A_{i,j}}{\sum_{j=1}^{N}A_{i,j}}$, where $N$ denotes the size of $S$. This allows the team leader to iteratively send invitations to preferred collaborators. Upon receiving an invitation, the invited scientist evaluates whether to join the team using the chain-of-thought process (Wei et al., 2022), considering the profiles of $s_{0}$ and the current team members. If accepted, the scientist is added to the team $T$. This process continues until the pre-defined team size $n$ is reached.
Topic Discussion. The next step is to propose a research topic, which will guide the research direction. Inspired by multi-round collaboration (Mezirow, 2003; Sunstein, 2005; Amgoud and Prade, 2009) and multi-agent collaboration strategies (Xu et al., 2023; Zhang et al., 2024; Shinn et al., 2024), we design an inter- and intra-team discussion mechanism. In this mechanism, team members engage in discussions based on a specific task description prompt (inter-team discussion). The default discussion mechanism between agents follows a round-table format with a sequential progression (Additional discussion topologies are presented in Appx. D.3.1). This process is also applied to subsequent collaboration steps. While allowing agents to decide when to stop the discussion would better reflect real-world scenarios, fixing the number of turns ensures consistent inference costs across different team settings in our experiments. Therefore, we leave the discussion of adaptive turn numbers to the ablation study (See Appx. D.3.2). The prompt for agent $i$ during the topic discussion is:
@@
Idea Generation. Third, the team is tasked with proposing several potential ideas. To align with genuine research workflows and mitigate LLM illusions (Huang et al., 2023), each agent is required to generate a comprehensive response that includes three key components: (1) a description of the idea, (2) a specific experimental plan, and (3) a self-assessment covering metrics such as novelty, feasibility, and clarity, representing the agent’s confidence (See Appx. Fig. 20).
@@
Inspired by the concept of gradually expanding an archive of ideas (Zhang et al., 2024; Lu et al., 2024), when a scientist $s_{i}$ at turn $k$ receives an existing idea from the previous response $R_{k,i-1}$, we retain the previously generated ideas along with their corresponding references from $B_{past}$. These are passed to the next agent, who can either refine the existing idea or propose a new one, depending on its choice. The prompt $Q_{k,i}$ is represented as:
@@
Afterwards, the response of $S_{i}$ at turn $k$ can be represented as $R_{k,i}\sim P_{s_{i}}(\cdot|Q_{k,i})$. After $K$ turns of discussion, we retain the three ideas with the highest confidence and store them in the idea list $I$.
@@
where $I_{j}$ is the $j$-th idea in $I$. Following chain-of-thought process, the response $R_{k,i}\sim P_{s_{i}}(\cdot|Q_{k,i})$ includes the scientist’s preferred idea and the reasoning behind their choice. The idea receiving the highest votes is then selected as the final idea, $R_{idea}$, for abstract generation. The effectiveness of novelty assessment is discussed in Appx. D.2.
Abstract Generation. Lastly, the team is required to produce a comprehensive abstract that includes the following sections: (1) Introduction, (2) Objective, (3) Methods, (4) Expected Results, and (5) Conclusion (Alexandrov and Hennerici, 2007). At the start of abstract generation, the team leader provides an initial draft based on $R_{idea}$. The first abstract-generation prompt is: $Q_{1,0}=\langle Q_{abstract},R_{idea}\rangle$, where $Q_{abstract}$ represents the task description and format requirements.
When an abstract is provided by the previous response $R_{k,i-1}$, the next scientist’s response should include: (1) an evaluation of the prior abstract (evaluation metrics are detailed in Appx. Fig. 23), (2) proposed modifications, and (3) the revised abstract to enable continuous refinement. The prompt is:
@@
A self-review mechanism is also considered after $R_{abstract}$ is finalized to pre-check its novelty. The optimized abstract $R_{abstract}$ is provided to the team leader to assess novelty by comparing it to similar papers in $B_{past}$ (See Appx. B.1 for more details). Because it introduces uncertainty in total inference cost, making it difficult to ensure fair experimental comparisons, we only discuss the effectiveness of this module in the ablation study (See Appx. D.2).